OpenStudy (anonymous):

The product of 3 and a number is 12 more than the sum of the number and 26. What is the number?

OpenStudy (midhun.madhu1987):

Let the number be x. 3x = (x + 26) + 12 Solve for x.

OpenStudy (anonymous):

is it 12

OpenStudy (anonymous):

@midhun.madhu1987

OpenStudy (midhun.madhu1987):

3x = (x + 26) + 12 3x - x = 26 + 12 2x = 38 x = ?

OpenStudy (anonymous):

19?

OpenStudy (midhun.madhu1987):

yeah :)

OpenStudy (anonymous):

thankyou i already gave you a medal

OpenStudy (midhun.madhu1987):

No Problem.. :)